# Reasoning Language Models: A Blueprint

### v3

@@ 11 Conclusion
The work also explores new directions in value and reward modeling, introducing a hierarchy of models and formally identifying several recent designs as instances of a new class of models, namely the Outcome-Driven Process Reward Model. This model class bridges the gap between outcome-based evaluation and process-based supervision by dynamically connecting intermediate reasoning steps to terminal outcomes, enabling more granular feedback during training without the need.
@@ Appendix A Mathematical Foundation of Markov Decision Processes for Reasoning Tasks
y so far to put this on the first page of the appendix
TABLE II: Overview of mathematical notation used in the paper
@@
$p(s^{\prime}\mid s,a)$ The probability of transition to state $s^{\prime}$ from state $s$ taking action $a$ in state $s$.
@@ A.1 Markov Decision Process
Markov Decision Process (MDP) is defined as a 5-tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space with $\mathcal{A}_{s}\subseteq\mathcal{A}$ denoting the set of actions which can be taken in the state $s$, $p$ represents the dynamics of transitions between states, i.e., $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ where $p(s,a,s^{\prime})$ is the probability of transitioning to the state $s^{\prime}$ when action $a$ was selected in the state $s$, $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function, i.e., $r(s,a,s^{\prime})$ represents the reward for arriving in the state $s^{\prime}$ after selecting the action $a$ in the state $s$, and $\gamma\in[0,1]$ is a discount factor.
@@ A.1.1 Solving an MDP
State value function $V_{\pi}(s_{t})$ represents the expected cumulative future reward for a given state $s_{t}$ under policy $\pi$:
$V_{\pi}(s_{t})=\mathbb{E}\left[\sum_{k=t}^{T}\gamma^{k-t}r(s_{k},a_{k},s_{k+1})\mid s_{t}\right]$ (3)
@@
The State-Action value function $Q(s_{t},a_{t})$ Oftentimes, it is useful to use the state-action value function $Q(s_{t},a_{t})$ instead of the state value function. Specifically, the state-action value function $Q(s_{t},a_{t})$ extends the state value function so that the function value is defined on a state and a specific action $a_{t}$:
@@ A.1.2 MDPs in the RLM Setting
$p(s,a,s^{\prime})=\begin{cases}1&\text{ if }z_{n+1}=z^{(a)}\\ 0&\text{ if }z_{n+1}\neq z^{(a)}\\ \end{cases}$
The definition of the reward function depends on the specific task. A reward commonly seen in reasoning tasks assigns non-zero reward only in the terminal states and hence only at the final reasoning step. This approach reflects the fact that for most tasks, the only final answer can be evaluated against the ground-truth solution to the original query. We call such reward functions sparse to clearly distinguish it from other setting in which intermediate rewards can be observed by the algorithm in the non-terminal states. The discount factor $\gamma$ determines how future rewards influence the current decision-making process. A higher discount factor ($\gamma\rightarrow 1$) places greater emphasis on long-term reasoning success, allowing the model to generate long reasoning sequences, while a lower discount factor prioritizes immediate rewards, incentivizing faster progress and shorter reasoning sequences.
@@
A.2 Monte Carlo Tree Search (MCTS)
@@
Edge statistics The MCTS algorithm stores the following three values for every edge $s,a$ in the search tree:
@@
  • $r(s,a)=r(s,a,s^{\prime})$ the reward received after taking the action $a$ in the state $s$ leading to the state $s^{\prime}$,
  • $\beta(s,a)$ the terminality function indicating if the action $a$ leads to a terminal state.
The Algorithm At the high level, the MCTS begins by initializing the tree with a single starting state $s_{0}$ as a root node and performing the following three phases in a loop:
@@
    Selection - a leaf-node in the current tree is selected for expanding its child (children).
@@ Appendix B Value and Reward Models
We now proceed to discuss details of value and reward models.
B.1 Outcome-Based Reward Models (ORM)
vs. Process-Based Reward Models (PRM)
@@ B.1 Outcome-Based Reward Models (ORM)
In reinforcement learning environments, reward models estimate the reward for taking an action $a$ in state $s$ which leads to state $s^{\prime}$. For reasoning tasks and algorithms like MCTS, which rely on evaluating intermediate steps, it is essential to have models capable of estimating the quality of each step. Two primary families of reward models for such process-based tasks are Outcome-Based Reward Models (ORMs) and Process-Based Reward Models (PRMs). Figure 12 compares both classes of models.
@@
Outcome-Based Reward Models (ORMs), first introduced by Uesato et al. [151], evaluate the reasoning process solely based on the final outcome. These models estimate the reward of the final step in the chain, often modeled in the literature as the likelihood of a correct final answer given the entire reasoning chain $P(correct(z_{T+1})\mid z_{0},...,z_{T+1})$ [151, 88] where $s_{T+1}:=z_{0},...,z_{T+1}$ is the complete reasoning chain consisting of reasoning steps $z_{i}$ and $T+1$ marks the last reasoning step. ORMs are particularly ill-suited for evaluating intermediate steps for several reasons. First, the training data and objective are inherently misaligned with step-wise evaluation, as they focus exclusively on final outcomes. Second, ORM evaluations tend to be overly pessimistic for intermediate steps since a subsequent erroneous step can obscure the correctness of earlier steps. This observation aligns with Havrilla et al. [59], who noted that ORMs often underestimate the solvability of a problem from an intermediate state and are prone to a high false-negative rate. Furthermore, ORMs lack robustness against false positives, potentially favoring erroneous reasoning steps and misleading the evaluation process.
@@
Process-Based Reward Models (PRMs), introduced by Lightman et al. [88] and Uesato et al. [151], evaluate reasoning on a step-by-step basis. These models estimate the reward of a step, which can be seen as the likelihood of correctness for the $t$-th step given its preceding context $P(correct(z_{t})\mid z_{0},...,z_{t})$ where $s_{t}:=z_{0},...,z_{t}$ is a potentially incomplete reasoning chain and $z_{i}$ are reasoning steps and $z_{0}$ is the query. PRMs provide more fine-grained feedback and can pinpoint errors in the chain. This step-wise evaluation provides dense rewards given partial responses and helps identify where reasoning deviates from correctness, offering improved interpretability and enabling more targeted improvements in reasoning processes. However, PRMs are computationally expensive to train and require extensive annotations of reasoning steps. These annotations, whether provided by humans or other LLMs, often suffer from limitations: human annotations are scarce, costly, and prone to bias, while prompted LLM-generated annotations [154] are typically of lower quality due to their limited self-evaluation capabilities [99]. Automated methods using for example MCTS such as [155, 96] introduce large computational costs and are prone to false negatives.
@@ B.2 Outcome-Driven Process-Based Reward Models
Motivated by the need for process-based reward models but constrained by the lack of annotated step-wise labels, certain models that we will refer to as Outcome-Driven Process-Based Reward Models (O-PRMs) have been proposed; they combine outcome-based signals with process-based objectives. We show these models in Figure 12. These models rely on process-based data, often automatically generated using MCTS algorithms, where simulations starting from a given step $s_{t}$ are performed. The final correctness of these simulated paths is aggregated to create step-wise labels [155, 96] (for other, non-MCTS approaches see [59]). This automation enables scalable data generation for O-PRMs, eliminating the need for extensive human annotation. Although O-PRMs can be categorized as process-based models due to their approximation of step-wise rewards, they remain inherently tied to outcome signals. Some authors [151] suggest that, under certain conditions, outcome signals in mathematical domains can approximate intermediate labels. However, O-PRMs inherit many limitations of ORMs, including susceptibility to false negatives, false positives, and an over-reliance on terminal outcomes. While the aggregation of multiple simulations helps reduce variance, the backtracking process may still oversimplify complex dependencies within reasoning chains.
Figure 12: Comparison of Outcome vs. Process-Based label generation, and the introduction of Outcome-Driven Process Based Reward Models (O-PRMs). Gray nodes mark terminal nodes.
@@
While the distinction between reward models and value models is often blurred in the literature—and their terminology is sometimes used interchangeably—we explicitly differentiate between these model types for evaluating reasoning steps. Additionally, we distinguish two variants of value modes: v-value and q-value models. This differentiation arises from the distinct roles these models play in reinforcement learning environments.
B.3.1 Reward Model (RM)
@@
A reward model predicts immediate rewards. In RL, this corresponds to the reward obtained for a transition $(s,a,s^{\prime})$ from state $s$ when taking action $a$ which results in step $s^{\prime}$. For reasoning, this corresponds to adding a new reasoning step $a$ to the structure. The new structure is then represented by $s^{\prime}$. Specifically, PRMs – which are preferred over ORMs for MCTS due to the need for action-based evaluation – learn these rewards and can be used to evaluate states (or the transition into a state). This formulation provides a localized, step-level evaluation independent of the overall outcome of the reasoning chain. The reward model is typically trained using labeled data where individual reasoning steps are associated with reward values. While this localized view is advantageous for step-by-step evaluation, it lacks the ability to consider how the current step contributes to the long-term success of the reasoning process. This limitation motivates the introduction of value models.
B.3.2 Value Model (VM)
@@
Value models provide a more abstract, global evaluation of states and actions by estimating their contribution to future rewards. Unlike reward models, which focus on immediate outcomes, value models consider both current and future rewards, enabling a broader perspective on reasoning quality. For example in reinforcement learning and MCTS, value models play a critical role in guiding the search process. By providing estimates of state or state-action values, they enable more informed decisions about which nodes to expand and explore. We now discuss variants of value models.
@@
V-Value Model (V-VM). One variant of a value model is the v-value model which predicts the expected cumulative future reward of a state, denoted as $V(s)$. This is equivalent to the state value function in reinforcement learning, which evaluates the long-term potential of the current state $s$. A key advantage of V-VMs is their global perspective, as they aggregate future rewards across all possible trajectories originating from the current state. However, V-VMs do not explicitly evaluate individual actions, which may limit their utility in step-level decision-making. Additionally, v-values are often ill-defined at terminal states, where rewards may substitute for state values during training.
@@
Q-Value Model (Q-VM). Another variant of a value model is the q-value model. Q-VMs predicts the expected cumulative future reward of taking a specific action $a$ in a given state $s$, denoted as $Q(s,a)$. Unlike V-VMs, Q-VMs explicitly associate values with state-action pairs, offering a more granular evaluation. This granularity makes Q-VMs particularly useful for MCTS, where decisions about which edge (action) to expand at a given node (state) are critical. By directly evaluating actions, Q-VMs align naturally with the selection mechanisms in MCTS, guiding the search toward promising paths. Similar to V-VMs, Q-VMs can also be categorized as PQVMs (Process-based Q-Value Models), OQVMs (Outcome-based Q-Value Models), and O-PQVMs (Outcome-driven Process-based Q-Value Models).
@@
Figure 13: Comparison of reward, v-value and q-value models in a sparse reward setting (only terminal states receive non-zero rewards). Gray nodes mark terminal nodes. The reward model should predict the rewards for transitioning from one state to another which is 0 for non-terminal states and not providing information. V-VMs and Q-VMs however, predict a global value and are therefore informative for non-terminal states.
@@ B.3.3 Example: Solving a Mathematical Equation
  • Reward Model (RM): A process-based reward model (PRM) might assign a reward $r(s_{t},a_{t},s_{t+1})$ for the reasoning step $a_{t}=\text{"Substitute }y=\sqrt{1-x^{2}}\text{"}$ ”. This reward quantifies the quality of the resulting state $s_{t+1}$, independent of whether it leads to a correct solution. However, in sparse reward settings (only final steps receive a reward), this reward would be 00.
  • V-Value Model (V-VM): A V-VM estimates $V(s_{t})$, representing the expected cumulative reward for the entire expected solution process starting from $s_{t}$. For instance, if $s_{t}=(\text{"Start with }x^{2}+y^{2}=1\text{"})$ ” ), $V(s_{t})$ considers the long-term potential of all reasoning paths originating from this state.
  • Q-Value Model (Q-VM): A Q-VM evaluates $Q(s_{t},a_{t})$, predicting the cumulative reward of taking a specific action $a_{t}$ (e.g., substituting $y=\sqrt{1-x^{2}}$) in state $s_{t}$. This value directly informs whether the action $a_{t}$ is likely to lead to a high-quality solution, providing a more granular evaluation compared to the V-VM.

### v4

@@ 11 Conclusion
The work also explores new directions in value and reward modeling, introducing a hierarchy of models and formally identifying several recent designs as instances of a new class of models, namely the Outcome-Driven Process-Based Reward Model. This model class bridges the gap between outcome-based and process-based evaluation by dynamically connecting intermediate reasoning steps to terminal outcomes, enabling more granular feedback during training without the need for extensive labels.
@@ Appendix A Mathematical Foundation of Markov Decision Processes for Reasoning Tasks
TABLE II: Overview of mathematical notation used in the paper.
@@
$p(s^{\prime}\mid s,a)$ The probability of transitioning to state $s^{\prime}$ from state $s$ by taking action $a$.
@@ A.1 Markov Decision Process
A Markov Decision Process (MDP) is defined as a 5-tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space with $\mathcal{A}_{s}\subseteq\mathcal{A}$ denoting the set of actions which can be taken in the state $s$, $p$ represents the dynamics of transitions between states, i.e., $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ where $p(s,a,s^{\prime})$ is the probability of transitioning to state $s^{\prime}$ when action $a$ was selected in state $s$, $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function, i.e., $r(s,a,s^{\prime})$ represents the reward for arriving in state $s^{\prime}$ after selecting action $a$ in state $s$, and $\gamma\in[0,1]$ is a discount factor.
@@ A.1.1 Solving an MDP
A state value function $V_{\pi}(s_{t})$ represents the expected cumulative future reward for a given state $s_{t}$ under policy $\pi$:
$V_{\pi}(s_{t})=\mathbb{E}\left[\sum_{k=t}^{T}\gamma^{k-t}r(s_{k},a_{k},s_{k+1})\mid s_{t}\right],$
@@
Oftentimes, it is useful to use a state-action value function $Q(s_{t},a_{t})$ instead of a state value function. Specifically, the state-action value function $Q(s_{t},a_{t})$ extends the state value function so that the function value is defined on a state and a specific action $a_{t}$:
@@ A.1.2 MDPs in the RLM Setting
$p(s,a,s^{\prime})=\begin{cases}1&\text{ if }z_{n+1}=z^{(a)}\\ 0&\text{ if }z_{n+1}\neq z^{(a)}\\ \end{cases}.$
The definition of the reward function depends on the specific task. A reward commonly seen in reasoning tasks assigns non-zero reward only in the terminal states and hence only at the final reasoning step. This approach reflects the fact that for most tasks, only the final answer can be evaluated against the ground truth solution to the original query. We call such reward functions sparse to clearly distinguish it from other settings in which intermediate rewards can be observed by the algorithm in the non-terminal states. The discount factor $\gamma$ determines how future rewards influence the current decision-making process. A higher discount factor ($\gamma\rightarrow 1$) places greater emphasis on long-term reasoning success, allowing the model to generate long reasoning sequences, while a lower discount factor prioritizes immediate rewards, incentivizing faster progress and shorter reasoning sequences.
@@
A.2 Monte Carlo Tree Search
@@
Edge statistics The MCTS algorithm stores the following values for every edge $s,a$ in the search tree:
@@
  • $r(s,a)=r(s,a,s^{\prime})$ the reward received after taking action $a$ in state $s$ leading to state $s^{\prime}$,
  • $\beta(s,a)$ the terminality function indicating if action $a$ leads to a terminal state.
Algorithm At the high level, the MCTS begins by initializing the tree with a single starting state $s_{0}$ as a root node and performing the following three phases in a loop:
@@
    Selection - a leaf node in the current tree is selected for expanding its child (children).
@@ Appendix B Value and Reward Models
We now proceed to discuss the details of value and reward models.
B.1 Outcome- vs. Process-Based Reward Models
@@
In reinforcement learning environments, reward models estimate the reward for taking an action $a$ in state $s$ which leads to state $s^{\prime}$. For reasoning tasks and algorithms like MCTS, which rely on evaluating intermediate steps, it is essential to have models capable of estimating the quality of each step. Two primary families of reward models for such tasks are Outcome-Based Reward Models (ORMs) and Process-Based Reward Models (PRMs). Figure 12 compares both classes of models.
Figure 12: Comparison of outcome- vs. process-based label generation, and the introduction of Outcome-Driven Process-Based Reward Models (O-PRMs). Gray nodes mark terminal nodes.
@@
Outcome-Based Reward Models, first introduced by Uesato et al. [155], evaluate the reasoning process solely based on the final outcome. These models estimate the reward of the final step in the chain, often modeled in the literature as the likelihood of a correct final answer given the entire reasoning chain $P(correct(z_{T+1})\mid z_{0},...,z_{T+1})$ [155, 91] where $s_{T+1}:=z_{0},...,z_{T+1}$ is the complete reasoning chain consisting of reasoning steps $z_{i}$ and $T+1$ marks the last reasoning step. ORMs are particularly ill-suited for evaluating intermediate steps for several reasons. First, the training data and objective are inherently misaligned with step-wise evaluation, as they focus exclusively on final outcomes. Second, ORM evaluations tend to be overly pessimistic for intermediate steps since a subsequent erroneous step can obscure the correctness of earlier steps. This observation aligns with Havrilla et al. [61], who noted that ORMs often underestimate the solvability of a problem from an intermediate state and are prone to a high false-negative rate. Furthermore, ORMs lack robustness against false positives, potentially favoring erroneous reasoning steps and misleading the evaluation process.
@@
Process-Based Reward Models, introduced by Lightman et al. [91] and Uesato et al. [155], evaluate reasoning on a step-by-step basis. These models estimate the reward of a step, which can be seen as the likelihood of correctness for the $t$-th step given its preceding context $P(correct(z_{t})\mid z_{0},...,z_{t})$ where $s_{t}:=z_{0},...,z_{t}$ is a potentially incomplete reasoning chain and $z_{i}$ are reasoning steps and $z_{0}$ is the query. PRMs provide more fine-grained feedback and can pinpoint errors in the chain. This step-wise evaluation provides dense rewards given partial responses and helps identify where reasoning deviates from correctness, offering improved interpretability and enabling more targeted improvements in reasoning processes. However, PRMs are computationally expensive to train and require extensive annotations of reasoning steps. These annotations, whether provided by humans or other LLMs, often suffer from limitations: human annotations are scarce, costly, and prone to bias, while prompted LLM-generated annotations [159] are typically of lower quality due to their limited self-evaluation capabilities [103]. Automated methods using for example MCTS such as [160, 100] introduce large computational costs and are prone to false negatives.
@@ B.2 Outcome-Driven Process-Based Reward Models
Motivated by the need for process-based reward models but constrained by the lack of annotated step-wise labels, certain models that we will refer to as Outcome-Driven Process-Based Reward Models (O-PRMs) have been proposed; they combine outcome-based signals with process-based objectives. We also illustrate these models in Figure 12. These models rely on process-based data, often automatically generated using MCTS algorithms, where simulations starting from a given step $s_{t}$ are performed. The final correctness of these simulated paths is aggregated to create step-wise labels [160, 100] (for other, non-MCTS approaches see [61]). This automation enables scalable data generation for O-PRMs, eliminating the need for extensive human annotation. Although O-PRMs can be categorized as process-based models due to their approximation of step-wise rewards, they remain inherently tied to outcome signals. Some authors [155] suggest that, under certain conditions, outcome signals in mathematical domains can approximate intermediate labels. However, O-PRMs inherit many limitations of ORMs, including susceptibility to false negatives, false positives, and an over-reliance on terminal outcomes. While the aggregation of multiple simulations helps reduce variance, the backtracking process may still oversimplify complex dependencies within reasoning chains.
@@
While the distinction between reward models and value models is often blurred in the literature—and their terminology is sometimes used interchangeably—we explicitly differentiate between these model types for evaluating reasoning steps. Additionally, we distinguish two variants of value modes: v-value and q-value models. This differentiation arises from the distinct roles these models play in reinforcement learning environments. We provide an overview over the differences between reward and value models in Figure 13.
Figure 13: Comparison of reward, v-value and q-value models in a sparse reward setting (only terminal states receive non-zero rewards). Gray nodes mark terminal nodes. The reward model should predict the rewards for transitioning from one state to another which is 0 for non-terminal states. V-VMs and Q-VMs however, predict a global value and are therefore informative for non-terminal states.
B.3.1 Reward Models
@@
A Reward Model (RM) predicts immediate rewards. In RL, this corresponds to the reward obtained for a transition $(s,a,s^{\prime})$ from state $s$ when taking action $a$ which results in step $s^{\prime}$. For reasoning, this corresponds to adding a new reasoning step $a$ to the structure. The new structure is then represented by $s^{\prime}$. Specifically, PRMs – which are preferred over ORMs for MCTS due to the need for action-based evaluation – learn these rewards and can be used to evaluate states (or the transition into a state). This formulation provides a localized, step-level evaluation independent of the overall outcome of the reasoning chain. The reward model is typically trained using labeled data where individual reasoning steps are associated with reward values. While this localized view is advantageous for step-by-step evaluation, it lacks the ability to consider how the current step contributes to the long-term success of the reasoning process. This limitation motivates the introduction of value models.
B.3.2 Value Models
@@
Value Models (VMs) provide a more abstract, global evaluation of states and actions by estimating their contribution to future rewards. Unlike reward models, which focus on immediate outcomes, value models consider both current and future rewards, enabling a broader perspective on reasoning quality. For example in reinforcement learning and MCTS, value models play a critical role in guiding the search process. By providing estimates of state or state-action values, they enable more informed decisions about which nodes to expand and explore. We now discuss variants of value models.
@@
One such variant is the V-Value Model (V-VM), which predicts the expected cumulative future reward of a state, denoted as $V(s)$. This is equivalent to the state value function in reinforcement learning, which evaluates the long-term potential of the current state $s$. A key advantage of V-VMs is their global perspective, as they aggregate future rewards across all possible trajectories originating from the current state. However, V-VMs do not explicitly evaluate individual actions, which may limit their utility in step-level decision-making. Additionally, v-values are often ill-defined at terminal states, where rewards may substitute for state values during training.
@@
Q-Value Models (Q-VMs) are another variant, which predict instead the expected cumulative future reward of taking a specific action $a$ in a given state $s$, denoted as $Q(s,a)$. Unlike V-VMs, Q-VMs explicitly associate values with state-action pairs, offering a more granular evaluation. This granularity makes Q-VMs particularly useful for MCTS, where decisions about which edge (action) to expand at a given node (state) are critical. By directly evaluating actions, Q-VMs align naturally with the selection mechanisms in MCTS, guiding the search toward promising paths. Similar to V-VMs, Q-VMs can also be categorized as Process-Based Q-Value Models (PQVMs), Outcome-Based Q-Value Models (OQVMs), and Outcome-Driven Process-Based Q-Value Models (O-PQVMs).
@@ B.3.3 Example: Solving a Mathematical Equation
  • Reward Model: A process-based reward model might assign a reward $r(s_{t},a_{t},s_{t+1})$ for the reasoning step $a_{t}=\text{"Substitute }y=\sqrt{1-x^{2}}\text{"}$ ”. This reward quantifies the quality of the resulting state $s_{t+1}$, independent of whether it leads to a correct solution. However, in sparse reward settings (only final steps receive a reward), this reward would be 00.
  • V-Value Model: A V-VM estimates $V(s_{t})$, representing the expected cumulative reward for the entire expected solution process starting from $s_{t}$. For instance, if $s_{t}=(\text{"Start with }x^{2}+y^{2}=1\text{"})$ ” ), $V(s_{t})$ considers the long-term potential of all reasoning paths originating from this state.
  • Q-Value Model: A Q-VM evaluates $Q(s_{t},a_{t})$, predicting the cumulative reward of taking a specific action $a_{t}$ (e.g., substituting $y=\sqrt{1-x^{2}}$) in state $s_{t}$. This value directly informs whether the action $a_{t}$ is likely to lead to a high-quality solution, providing a more granular evaluation compared to the V-VM.